# Metadata practices for simulation workflows

### v1

@@ 3 Metadata post-processing framework – the Archivist
To illustrate the Archivist framework, we show a representative example implementation a workflow to parse and structure metadata in fig. 2. The framework, coded in Python (http://www.python.org, last access: 23 February 2024), is primarily composed of four processing classes and one interface class. The processing classes perform different aspects of the metadata parsing and structuring (exploration of metadata files, parsing of metadata sources, formatting and structuring of collected metadata, and exporting to a selected format), while the central interface class orchestrates the other classes for ease of use.
@@
Figure 2: Implementation of an Archivist metadata processing pipeline with two example parsers. From the user perspective, the Archivist class is provided with all inputs associated with the pipeline. Internally, the Explorer class sorts the individual files to process from a collection of raw metadata files, and dispatches them to corresponding Parsers classes, here the Configuration parser and the Time parser classes. Then, each Parser class employs a user defined function to extract specific information from its respective files. After this, the Formatter class collects the parsing results. If a schema is provided, the composite result can be structured following user design. The final processed metadata is output in a format of choice by the Exporter class. Bottom: Legend for different components of the figure.
@@
The Archivist class is a convenience interface class which instantiates and orchestrates the processing classes. As input, the class accepts a collection of raw metadata, the parsing operations to apply, and optionally a formatting schema. This interface is necessary because each processing class uses specifically structured inputs and outputs, which can be cumbersome and error-prone for users to generate themselves. We designed the interface to be configurable, giving it enough flexibility to customize the behavior of the processing classes while providing simple inputs and outputs.
Explorer:
@@
The Explorer class processes inputs given to the Archivist. These can be raw metadata archives or directories containing raw metadata files. To enhance interoperability, no assumption is made on the structure and contents of the archives or directories. As such the user must define rules to identify which files to parse. To do this, the user can provide precise file names or regular expressions describing these names. We refer to these as file description rules. Using these rules, the explorer searches the input (archive or directory) for corresponding files and provides a list of files to parse.
Parser:
The Parser is used to gather metadata from a subset of the list of files identified by the Explorer. For this, the user must associate one of the file description rules provided to the Explorer with a Parser instance. Depending on the rule, the Archivist instance dispatches each file to the respective Parser. The Parser is an abstract class designed to be extended by users for their own purposes. Such an extended instance contains a file description rule and an associated parsing method. Although each user must provide their own parsing methods, the extension mechanism allows transparent file parsing regardless of format or structure. Furthermore, with the source code, we provide examples of parsing methods we use for our own workflows and we hope to build a user base around shared methods to foster reusability and replicability.
Formatter:
@@
The Formatter combines the output of the Parser instances into a unified metadata file. Although parsing the desired metadata files and listing the results in a single file might be sufficient for some workflows, in other cases it may be necessary to transform the collection of parsing results into a cohesive and comprehensive structure. For this, the user must provide a schema to match a parsing result to the desired structure. We based the format of our schema on the JSON Schema standard[23], and defined additional directives to orchestrate more complex transformations. As a simplified example in fig. 2, given the parsed information from the configuration file and time file, the user can combine the simulation time with the real time in a single field, as the real_time_factor .
Exporter:
@@
The Exporter saves the internal data representation of the structured metadata to a file with specific output format. Like the Parser, though not an abstract class itself, this class was designed to be extended to enable serialization to arbitrary formats. In particular, users define the output format that enhances compatibility with their annotation method of choice. In fig. 2, the Archivist employs the JSON format to export the structured metadata.
@@
Using this class hierarchy, users only need to define the parsing functionalities they are interested in along with the file target rules, and provide a metadata collection to process. The Archivist class takes care of coordinating the other classes and generating the output. For more complex operations, users can provide a schema for the Formatter class, and extend the Exporter class to change the output format. With this flexibility, the Archivist framework provides a re-usable parsing and structuring pipeline that can be operated on existing metadata or attached to a workflow for automated post-processing.
4 Examples
@@
To illustrate how the proposed metadata practices (“Metadata management practices”) and the Archivist framework (“Metadata post-processing framework – the Archivist”) can be implemented into specific simulation workflows, we here provide a minimal and two real-world examples. The first example (“Minimal example”) constitutes a toy workflow applicable to generic simulations where the user conducts a parameter scanning experiment and leverages recorded metadata to identify suitable configurations. The second example (“Neuroscience use case”) describes a workflow for the benchmarking and the verification of a specific neuroscience simulation architecture. The third example (“Hydrology use case”) showcases a routine procedure for the calibration of a hydrological model.
4.1 Minimal example
@@
The first example mimics a typical problem in computational science (fig. 3): how to choose the parameters of a given model such that the accuracy of its predictions is high while the time-to-solution is small? In many applications, a typical parameter which improves the model prediction but slows down the simulation is the model size, here referred to as the scale. Examples are the number of elements in a finite element simulation, or the number of neurons in an artificial neuronal network. The minimal example described here illustrates how the enrichment of simulation results with processed metadata helps finding an answer to the above question. For simplicity, we assume that the example runs in a local simulation environment where the required software stack is already installed, no job manager is used, and the data is stored locally. In accordance with fig. 1, we subdivide the entire workflow into the three components: Simulation, Metadata post-processing, and Data usage. In the following, we describe each of these components in detail.
@@
Figure 3: Minimal example. Illustration of the Archivist’s functionality in a simple example use case. In a parameter scanning experiment, several instances of a model with different configurations (parameters) are simulated (“Simulation $1$”,…, “Simulation $N$”; blue boxes on the left). During each simulation, configuration and performance information are recorded and stored in a (raw) metadata archive (yellow). After each simulation, the stored metadata is post-processed (“Metadata post-processing $1$”,…, “Metadata post-processing $N$; red): first, the relevant information is extracted by calling a predefined parsing script (gray box parsers.py). Non-relevant information is discarded (see light gray text in the raw metadata files). The extracted metadata are then structured according to a provided schema (gray box schema.json). Finally, the simulation results are annotated with the structured metadata and stored in a database (red cylinder). After all simulations and the metadata post-processing are finished and their results stored in the database, the annotated data can be queried and presented (green).
@@
The Simulation section consists of the configuration of the model and the simulation architecture, the simulation execution, and the collection of the generated data and metadata (fig. 3, blue box). Here, we do not further specify the actual structure and dynamics of the underlying hypothetical model. We merely assume that it contains a parameter scale representing the model size. During the simulation, the model dynamics is propagated forwards in time with some time resolution (step_size), up to a total model time (sim_time). We further assume that the simulation architecture is equipped with a parallelization infrastructure, which is parameterized by the total number of processes (procs) and the number of computing threads used by each process (threads). Model and architecture parameters are stored in a configuration file (config.yaml). During each simulation, some primary simulation data (which we do not further specify here) is generated and stored in a data file (results.dat). In addition, the simulation scripts monitor the duration of each simulation run with the help of the linux time command (https://www.man7.org/linux/man-pages/man1/time.1.html, last access: 25 August 2024). It returns three different types of durations: the real time represents the actual wall-clock time, i.e., the total time taken by the simulation process from invocation to termination. The user and the sys times measure the cumulative CPU (thread) time spent in the user and in the kernel mode, respectively. All three times (real, user, sys) are stored in a text file (time.txt). The simulation is repeated for a range of scale’s, procs, and threads. At the end of each simulation run, the configuration and the time files are collected in the form of a raw metadata archive (metadata_archive.tgz).
@@
During the Metadata post-processing, relevant configuration information and simulation times are parsed from the raw metadata archive and combined into a structured metadata file (fig. 3, red box). For the question at hand, not all metadata collected during the simulation needs to be extracted. Let’s, for example, assume that the time resolution (step_size) was kept constant for all simulations. Moreover, we are only interested in the real but not in the user and in the sys times. During the parsing of the raw metadata, only a subset of metadata is therefore extracted (black items in config.yaml and time.txt in fig. 3). In addition, physical units are added to the time quantities. The metadata extraction is performed by the Archivist with the help of user-defined parsing methods (parsers.py). The Archivist further structures the extracted metadata according to a user-defined schema (schema.json). In this example, the schema introduces the total number of virtual processes (virtual_processes) as the product of the number of processes and the number of threads per process, as well as the real time factor (real_time_factor) as the ratio between the measured wall-clock (real) time and the model time (sim_time). Finally, the primary simulation data are annotated by the structured metadata. In this example, this annotation is performed by uploading the primary simulation data and the corresponding structured metadata as a single entry to a local data base. An implementation of the metadata post-processing performed for this example can be found at https://doi.org/10.5281/zenodo.13442425 (last access: 30 August 2024).
@@
In the Data usage section, the local data base containing the accumulated annotated simulation results is queried to ultimately find model sizes where the model accuracy is sufficiently high while the time-to-solution is low (fig. 3, green box). To this end, the parameter search is restricted to a subset of simulations where a given amount of computational ressources has been used. In this example, only those entries are extracted from the data base where the total number of virtual processes equals $16$. Based on the selected primary simulation data, the user assesses the model accuracy, for example, by comparing the model predictions with some observed or experimental data (for a concrete example, see section 4.3). At the same time, the user extracts the real time factor from the corresponding structured metadata. A subsequent analysis of the dependence of the model accuracy and the real time factor on the model size (scale), and an account of additional constraints such as the maximum acceptable real time factor or the minimal acceptable model accuracy, permits an identification of appropriate model sizes.
4.2 Neuroscience use case
@@
Understanding how the brain “computes”, what principles it employs to solve complex tasks with minimal energy consumption, how it evolves and changes during the lifetime of an organism, what the origins and effects of neurodegenerative diseases are and what possibilities of treatment exist has huge social, economical and ecological impact. The human brain consists of about $10^{11}$ nerve cells (neurons)[24], which form a complex network. Each neuron receives inputs from thousands of other neurons, both from the local neighborhood and from distant brain areas. The connectivity structure is highly heterogeneous and depends on the involved neuron types and brain areas. Furthermore, the connections between neurons (synapses) are not static but change depending on sensory inputs and other factors. The mathematical description of the brain’s dynamics at cellular resolution therefore involves large sets of coupled differential equations. Even for a single cubic millimeter of brain tissue, this number would be on the order of at least $10^{4}$. Neuroscience is thus dealing with complicated mesoscopic dynamical systems, which are neither small nor in the thermodynamic limit. They can not fully be understood by means of analytical mathematical methods from dynamical systems theory or statistical physics. State-of-the-art neuroscience hence relies on simulation.
@@
Simulating brain-scale neuronal networks at cellular resolution, i.e., instantiating the corresponding models and simulating them in a reasonable time, is challenging. In particular, the investigation of slow biological processes, such as learning or brain development, requires accelerated simulations where the wall-clock times $T_{\text{wall}}$ are substantially smaller than the duration $T_{\text{model}}$ of the simulated time interval. Due to the large number of neurons and connections, brain-scale neuronal-network simulation requires substantial amounts of memory to store all involved state variables[25]. At each instance of time, large numbers of differential equations have to be solved simultaneously. Brain simulation thus typically employs parallel, distributed computing. One of the key challenges in distributed brain simulation, however, is the communication between neurons, which are typically located on different compute nodes. The development of efficient algorithms and simulation software that can exploit the possibilities of continuously evolving high-performance computing architectures and incorporate new insights from experimental and theoretical neuroscience is hence a fundamental activity in this field[26, 27, 28, 29, 30, 31, 32, 33, 34]. It depends on large teams of software developers, who coordinate their work over years of development with many and fast update cycles. Continuously monitoring the quality (correctness, usability, reproducibility) and the performance (speed, memory demands, energy costs) of the simulation code is mandatory[3, 35]. Without regularly testing the simulation code, the integration of new features or new optimizations may unknowingly lead to wrong results or performance breakdowns[14]. Detecting and understanding such unwanted behavior, comparing different hardware configurations or testing procedures, and sharing test results with other developers relies on an efficient tracking and organization of the corresponding benchmarking data and metadata.
@@
The use case presented here demonstrates how the proposed metadata managing guidelines and the Archivist can help in fulfilling this task. It illustrates a verification and performance benchmarking workflow for the neuronal-network simulation code NEST GPU[36, 37] executed on a range of different hardware platforms (fig. 4). It shows that the simulation code generates identical results on all tested platforms (fig. 4B), and highlights differences in the simulation speed (fig. 4C). For illustration, we restrict this example to
@@
The entire pipeline underlying this use case is implemented using Snakemake[16] – a lightweight yet powerful and flexible workflow manager. Individual workflow components, such as the software deployment and compilation, the simulation execution, the storage of simulation data and raw metadata, the data and metadata post-processing, as well as the data exploration and presentation, correspond to specific Snakemake rules (for details, see “Methods”). During each execution of the simulation workflow, the (probabilistic) network model is simulated with $10$ different random realizations of the initial state and the network connectivity (different random-number-generator [RNG] seeds). Each simulation instance generates single-neuron activity traces (spike trains) as well as population and time averaged firing rates as the primary neuroscientific data (blue cylinder in fig. 4A). In addition, information about the model parameterization, RNG seeds, the hard- and software configuration, as well as the wall-clock times and real time factors are stored in various files and formats as raw metadata (yellow cylinder in fig. 4A). After the simulation, the data and raw metadata are compressed, labeled with a unique identifier (uid), and uploaded to an instance of a Mongo database (MongoDB, https://www.mongodb.com, last access: 23 February 2024). During the metadata post-processing (red gear in fig. 4A), the Archivist uses specific parsers to extract the type of the computing platform, the RNG seed, and the real time factor from the raw metadata archives. Subsequently, the extracted information is structured according to a user-specific schema. The link between the data and the metadata (data annotation) is established by attaching the data uid to the structured metadata. At the end of the metadata post-processing, the structured metadata are uploaded to the database containing the data and raw metadata (red cylinder in fig. 4A). The workflow is executed on four different GPU platforms, each sweep individually enriching the same database with its raw and structured metadata and annotated data. The database containing the accumulated data and metadata from various instantiations of the simulation and metadata post-processing workflow can be efficiently queried according to user interest (curved red arrows in fig. 4A). For the verification and the performance assessment of the NEST GPU code, we extract the average firing rates and the real time factors, respectively, for each network realization (RNG seed) and computing platform, and plot the corresponding across-trial averages and standard deviations (fig. 4B,C).
@@
Figure 4: Neuroscience use case. A) Left: Sketch of a model [38] describing the activity dynamics generated by a local neuronal circuit of the mammalian neocortex (adapted from [40]). The network model is composed of four excitatory (E; blue triangles) and four inhibitory neuronal populations (I; red circles), distributed across four cortical layers 2/3, 4, 5 and 6, and driven by background inputs. Neurons in the network are interconnected in a cell-type and layer specific manner (blue and red arrows). Right: The model generates neuronal activity data (“spikes” and firing rates) as the primary neuroscientific data (blue cylinder). For each simulation instance, information about the model parameterization, random number generator (RNG) seeds, the hard- and software configuration, as well as the wall-clock times are stored in various files and formats as raw metadata (yellow cylinder). In a subsequent post-processing step (red gear), the metadata is parsed and structured by the Archivist. The simulation data is annotated with this structured metadata and stored in a database for further usage (red cylinder). The database can flexibly be queried according to user interests (curved red arrows). B, C) Verification (B) and performance benchmarking (C) as two exemplary types of data usage. B) Average activity level (firing rate) in each of the 8 neuronal populations 2/3E, …, 6I depicted in panel A, obtained from simulations of the model on four different GPU platforms (see labels at horizontal axis in panel C). C) Real time factor (ratio between wall-clock time $T_{\text{wall}}$ and simulated biological time $T_{\text{model}}=10\,\text{s}$) for four different GPU computing platforms. Error bars (red) in B) and C) depict standard deviations across ten different model realizations (random-number generator [RNG] seeds) and simulation runs for each platform (error bars are partly too small to be visible).
4.3 Hydrology use case
@@
The simulation of the hydrologic cycle is fundamental in environmental modeling. The movement and storage of water in the terrestrial system includes diverse processes. Key processes are the infiltration of precipitation into the subsurface, the storage of water in soils and the removal of water from the soil via plant transpiration and subsurface runoff. A reliable estimation of water fluxes and storages is relevant for a wide range of sectors, like drinking water supply, agriculture, forestry, energy production, and transportation. Hydrologic models simulate hydrologic state variables and fluxes and their interaction. These models are created for different purposes according to the needs of specific sectors outlined above. These range from infrastructure planning for drinking water supply to drought quantification for agriculture and to climate projections for future water management. Hydrologic models represent fundamental processes like snow accumulation and melting, soil infiltration, evaporation and plant transpiration, surface and subsurface runoff, and river routing. Typical output variables are river discharge, evapotranspiration and soil moisture among others. Input variables are precipitation and air temperature. There are a wide range of computational workflows given the diverse purposes that hydrologic models are used for. One of the most complex workflows are related to climate change projections where input variables are taken from different sources like climate model ensembles[44]. Currently, metadata tracking in such workflows is often very limited and does not contain the comprehensive settings that are necessary to execute these workflows.

### v2

@@ Results
To illustrate the Archivist framework, we show a representative example implementation of a workflow to parse and structure metadata in fig. 2. The framework, coded in Python (http://www.python.org, last access: 23 February 2024), is primarily composed of four processing classes and one interface class. The processing classes perform different aspects of the metadata parsing and structuring (exploration of metadata files, parsing of metadata sources, formatting and structuring of collected metadata, and exporting to a selected format), while the central interface class orchestrates the other classes .
@@
Figure 2: Implementation of an Archivist metadata processing pipeline with two example parsers. From the user perspective, the Archivist class is provided with all inputs associated with the pipeline (bottom yellow and grey boxes). Internally, the Explorer class extracts the individual files to process from a collection of raw metadata files, and dispatches them to corresponding Parser classes, here the Configuration parser and the Time parser classes. Then, each Parser class employs a user-defined function to extract specific information from its respective files. After this, the Formatter class collects the parsing results. If a data template is provided, the composite result can also be restructured. The final processed metadata is output in a format of choice by the Exporter class. A Jupyter tutorial explaining how to implement the metadata processing pipeline shown here is available at https://github.com/INM-6/metadata-archivist/blob/main/examples/schema_example4/schema_tutorial.ipynb (last access: 26 January 2025).
@@
The Archivist class is a convenience interface class which instantiates and orchestrates the processing classes. As input, the class accepts a collection of raw metadata, the parsing operations to apply, and optionally a data template. This interface is necessary because each processing class uses specifically structured inputs and outputs, which can be cumbersome and error-prone for users to generate themselves. We designed the interface to be configurable, giving it enough flexibility to customize the behavior of the processing classes while providing simple inputs and outputs.
Explorer class:
@@
The Explorer class processes inputs given to the Archivist. These can be raw metadata archives or directories containing raw metadata files. To enhance flexibility, no assumption is made about the structure and contents of the archives or directories. As such, the user must define rules identifying which files to parse – we refer to this as a file target rule. To do this, the user can provide precise file names or regular expressions describing these names. We refer to these as file description rules. Using these rules, the Explorer searches the input (archive or directory) for corresponding files and provides a list of files to parse.
Parser abstract class:
The Parser is an abstract classed designed to be extended by users to extract metadata from files. When instantiating a Parser, the user must associate one of the file target rules provided to the Explorer . Depending on the rule, the Archivist instance dispatches each file to the respective Parser.
Additionally each user must provide their own parsing methods. Two examples of such instances are shown in fig. 2: the config_parser which employs the PyYAML (https://github.com/yaml/pyyaml, last visited: 26 January 2025) package to read YAML files, and the time_parser which uses a custom ASCII file reader. The ASCII reader consists of a simple line-by-line parsing loop over the input file, using a regular expression to select only lines starting with a word followed by a space then a time marker composed of minutes and seconds. If the word matches the string “real", then the time marker is stored and its time value converted to seconds. The final output is a dictionary containing the value in seconds of the real time and a unit description. Additional examples of parsing classes are provided with our source code and we hope to build a user base around shared methods to foster reusability and replicability.
Formatter class:
@@
The Formatter combines the output of the Parser instances into a unified metadata file. Although parsing the desired metadata files and listing the results in a single file might be sufficient for some workflows, in other cases it may be necessary to transform the collection of parsing results into a cohesive and comprehensive structure. For this, the user must provide a data template to match a parsing result to the desired structure. These data templates are defined with an extended implementation of the JSON Schema[25] which the Formatter can interpret to link data structures with extracted information. As a simplified example in fig. 2, given the parsed information from the configuration file and time file, the user can combine the simulation time with the real time in a single field as the real_time_factor .
Exporter class:
@@
The Exporter saves the internal data representation of the structured metadata to a file with specific output format. Like the Parser, though not an abstract class itself, this class was designed to be extensible to enable serialization to arbitrary formats. In particular, users define the output format that enhances compatibility with their annotation method of choice. In fig. 2, the Archivist employs the JSON format to export the structured metadata.
@@
Using this class hierarchy, users only need to define the parsing functionalities they are interested in along with the file target rules, and provide a metadata collection to process. The Archivist class takes care of coordinating the other classes and generating the output. For more complex operations, users can provide a data template for the Formatter class, and extend the Exporter class to change the output format. With this flexibility, the Archivist framework provides a re-usable parsing and structuring pipeline that can operate on existing metadata or be attached to a workflow for automated post-processing. A Jupyter tutorial explaining how to implement the metadata processing pipeline shown in fig. 2 is available at https://github.com/INM-6/metadata-archivist/blob/main/examples/schema_example4/schema_tutorial.ipynb (last access: 26 January 2025).
Examples
@@
To illustrate how the proposed metadata practices (“Metadata management practices”) and the Archivist framework (“Metadata post-processing framework – the Archivist”) can be implemented into specific simulation workflows, we provide here one minimal and two real-world examples. The first example (“Minimal example”) constitutes a toy workflow applicable to simulations where the user conducts a parameter scanning experiment and leverages recorded metadata to identify suitable configurations. The second example (“Neuroscience use case”) describes a workflow for the benchmarking and the verification of a specific neuroscience simulation architecture. The third example (“Hydrology use case”) showcases a routine procedure for the calibration of a hydrological model.
Minimal example
@@
The first example addresses a typical question in computational science (fig. 3): how can we choose the parameters of a given model such that the accuracy of its predictions is high while the time-to-solution is small? In many applications, a typical parameter which improves the model prediction but slows down the simulation is the model size, here referred to as the scale. Examples are the number of elements in a finite element simulation, or the number of neurons in an artificial neuronal network. The minimal example described here illustrates how the enrichment of simulation results with processed metadata helps find an answer to the above question. For simplicity, we assume that the example runs in a local simulation environment where the required software stack is already installed, no job manager is used, and the data is stored locally. In accordance with fig. 1, we subdivide the entire workflow into the three components: Simulation, Metadata post-processing, and Data usage. Below, we describe each of these components in detail.
@@
Figure 3: Minimal example. Illustration of the Archivist’s functionality in a simple example use case. In a parameter scanning experiment, several instances of a model with different configurations (parameters) are simulated (“Simulation $1$”,…, “Simulation $N$”; blue boxes on the left). During each simulation, configuration and performance information are recorded and stored in a (raw) metadata archive (yellow). After each simulation, the stored metadata is post-processed (“Metadata post-processing $1$”,…, “Metadata post-processing $N$”; red): first, the relevant information is extracted by user-defined Parser classes (gray box parsers.py). Non-relevant information is discarded (see light gray text in the raw metadata files). The extracted metadata are then structured according to a provided data template (gray box data_template.json). Finally, the simulation results are annotated with the structured metadata and stored in a database (red cylinder). After all simulation and metadata post-processing instances are finished and their corresponding results are stored in the database, the annotated data can be queried and presented (green).
@@
The Simulation section consists of the configuration of the model and the simulation architecture, the simulation execution, and the collection of the generated data and metadata (fig. 3, blue box). Here, we do not further specify the actual structure and dynamics of the underlying hypothetical model. We merely assume that it contains a parameter scale representing the model size. During the simulation, the model dynamics are propagated forward in time with some time resolution (step_size), up to a total model time (sim_time). We further assume that the simulation architecture is equipped with a parallelization infrastructure, which is parameterized by the total number of processes (procs) and the number of computing threads used by each process (threads). Model and architecture parameters are stored in a configuration file (config.yaml). During each simulation, some primary simulation data (which we do not further specify here) is generated and stored in a data file (results.dat). In addition, the simulation scripts monitor the duration of each simulation run with the help of the Linux time command (https://www.man7.org/linux/man-pages/man1/time.1.html, last access: 25 August 2024). It returns three different types of durations: the real time represents the actual wall-clock time, i.e., the total time taken by the simulation process from invocation to termination, while the user and the sys times measure the cumulative CPU (thread) time spent in the user and in the kernel mode, respectively. All three times (real, user, sys) are stored in a text file (time.txt). The simulation is re-run for a range of scale’s, procs, and threads. At the end of each simulation run, the configuration and the time files are collected in the form of a raw metadata archive (metadata_archive.tgz).
@@
Within the Metadata post-processing section, relevant configuration information and simulation times are parsed from the raw metadata archive and combined into a structured metadata file (fig. 3, red box). For the question at hand, not all metadata collected during the simulation needs to be extracted. Let’s, for example, assume that the time resolution (step_size) was kept constant for all simulations. Moreover, we are only interested in the real but not in the user and in the sys times. During the parsing of the raw metadata, only a subset of metadata is therefore extracted (black items in config.yaml and time.txt in fig. 3). In addition, physical units are added to the time quantities. The metadata extraction is performed by the Archivist with the help of user-defined parsing classes (parsers.py). The Archivist further structures the extracted metadata according to a user-defined template (data_template.json). In this example, the template introduces the total number of virtual processes (virtual_processes) calculated by multiplying the number of processes by the number of threads per process, as well as the real time factor (real_time_factor) calculated by taking the ratio between the measured wall-clock (real) time and the model time (sim_time). Finally, the primary simulation data are annotated by the structured metadata. In this example, annotation is performed by uploading the primary simulation data and the corresponding structured metadata as a single entry to a local database. An implementation of the metadata post-processing performed for this example can be found at https://doi.org/10.5281/zenodo.13442425 (last access: 30 August 2024).
@@
In the Data usage section, the local database containing the accumulated annotated simulation results is queried to ultimately find model sizes where the model accuracy is sufficiently high while the time-to-solution is low (fig. 3, green box). To this end, the parameter search is restricted to a subset of simulations where a given amount of computational ressources have been used. In this example, only those entries are extracted from the database where the total number of virtual processes equals $16$. Based on the selected primary simulation data, the user assesses the model accuracy, for example, by comparing the model predictions with some observed or experimental data (for a real-world example, see “Hydrology use case”). At the same time, the user extracts the real time factor from the corresponding structured metadata. A subsequent analysis of the dependence of the model accuracy and the real time factor on the model size (scale), and an account of additional constraints such as the maximum acceptable real time factor or the minimal acceptable model accuracy, permits an identification of appropriate model sizes.
Neuroscience use case
@@
Understanding how the brain “computes”, what principles it employs to solve complex tasks with minimal energy consumption, how it evolves and changes during the lifetime of an organism, what the origins and effects of neurodegenerative diseases are, and what possibilities of treatment exist has huge social, economical, and ecological impact. The human brain consists of about $10^{11}$ nerve cells (neurons)[26], which form a complex network. Each neuron receives inputs from thousands of other neurons, both from its local neighborhood and from distant brain areas. The connectivity structure is highly heterogeneous and depends on the involved neuron types and brain areas. Furthermore, the connections between neurons (synapses) are not static but change depending on sensory inputs and other factors. The mathematical description of the brain’s dynamics at cellular resolution therefore involves large sets of coupled differential equations. Even for a single cubic millimeter of brain tissue, this number is on the order of at least $10^{4}$. Neuroscience is thus dealing with complicated mesoscopic dynamical systems, which are neither small nor in the thermodynamic limit. They can not be fully understood by means of analytical mathematical methods from dynamical systems theory or statistical physics. State-of-the-art neuroscience hence relies on simulation.
@@
Simulating brain-scale neuronal networks at cellular resolution, i.e., instantiating the corresponding models and simulating them in a reasonable time, is challenging. In particular, investigating slow biological processes, such as learning or brain development, requires accelerated simulations where the wall-clock times $T_{\text{wall}}$ are substantially smaller than the duration $T_{\text{model}}$ of the simulated time interval. Due to the large number of neurons and connections, brain-scale neuronal-network simulation requires substantial amounts of memory to store all involved state variables[27]. At each instance of time, large numbers of differential equations have to be solved simultaneously. Brain simulation thus typically employs parallel, distributed computing. One of the key challenges in distributed brain simulation, however, is the communication between neurons, which are typically located on different compute nodes. Developing efficient algorithms and simulation software that can exploit the possibilities of continuously evolving high-performance computing architectures and incorporate new insights from experimental and theoretical neuroscience is hence a fundamental activity in this field[28, 29, 30, 31, 32, 33, 34, 35, 36]. It depends on large teams of software developers, who coordinate their work over years of development with many and fast update cycles. Continuously monitoring the quality (correctness, usability, reproducibility) and the performance (speed, memory demands, energy costs) of the simulation code is mandatory[3, 37]. Without regularly testing the simulation code, the integration of new features or new optimizations may unknowingly lead to wrong results or performance breakdowns[14]. Detecting and understanding such unwanted behavior, comparing different hardware configurations or testing procedures, and sharing test results with other developers relies on an efficient tracking and organization of the corresponding benchmarking data and metadata.
@@
The use case presented here demonstrates how the proposed metadata management practices and the Archivist can help in fulfilling this task. It illustrates a verification and performance benchmarking workflow for the neuronal-network simulation code NEST GPU[38, 39] executed on a range of different hardware platforms (fig. 4). It shows that the simulation code generates identical results on all tested platforms (fig. 4B), and highlights differences in the simulation speed (fig. 4C). For illustration, we restrict this example to
@@
The entire pipeline underlying this use case is implemented using Snakemake[16] – a lightweight yet powerful and flexible workflow manager. Individual workflow components, such as the software deployment and compilation, the simulation execution, the storage of simulation data and raw metadata, the data and metadata post-processing, as well as the data exploration and presentation, correspond to specific Snakemake rules (for details, see “Methods”). During each execution of the simulation workflow, the (probabilistic) network model is simulated with $10$ different random realizations of the initial state and the network connectivity (different random-number-generator [RNG] seeds). Each simulation instance generates single-neuron activity traces (spike trains) as well as population and time averaged firing rates as the primary neuroscientific data (blue cylinder in fig. 4A). In addition, information about the model parameterization, RNG seeds, the hard- and software configuration, as well as the wall-clock times and real time factors are stored in various files and formats as raw metadata (yellow cylinder in fig. 4A). After the simulation, the data and raw metadata are compressed, labeled with a unique identifier (uid), and uploaded to an instance of a Mongo database (MongoDB, https://www.mongodb.com, last access: 23 February 2024). During the metadata post-processing (red gear in fig. 4A), the Archivist uses specific parsers to extract the type of the computing platform, the RNG seed, and the real time factor from the raw metadata archives. Subsequently, the extracted information is structured according to a user-specific template. The link between the data and the metadata (data annotation) is established by attaching the data uid to the structured metadata. At the end of the metadata post-processing, the structured metadata are uploaded to the database containing the data and raw metadata (red cylinder in fig. 4A). The workflow is executed on four different GPU platforms, each sweep individually enriching the same database with its raw and structured metadata and annotated data. The database containing the accumulated data and metadata from various instantiations of the simulation and metadata post-processing workflow can be efficiently queried according to user interest (curved red arrows in fig. 4A). For the verification and the performance assessment of the NEST GPU code, we extract the average firing rates and the real time factors, respectively, for each network realization (RNG seed) and computing platform, and plot the corresponding across-trial averages and standard deviations (fig. 4B,C).
@@
Figure 4: Neuroscience use case. A) Left: Sketch of a model [40] describing the activity dynamics generated by a local neuronal circuit of the mammalian neocortex (adapted from [42]). The network model is composed of four excitatory (E; blue triangles) and four inhibitory neuronal populations (I; red circles); distributed across four cortical layers 2/3, 4, 5 and 6; and driven by background inputs. Neurons in the network are interconnected in a cell-type and layer specific manner (blue and red arrows). Right: The model generates neuronal activity data (“spikes” and firing rates) as the primary neuroscientific data (blue cylinder). For each simulation instance, information about the model parameterization, random number generator (RNG) seeds, the hard- and software configuration, as well as the wall-clock times are stored in various files and formats as raw metadata (yellow cylinder). In a subsequent post-processing step (red gear), the metadata is parsed and structured by the Archivist. The simulation data is annotated with this structured metadata and stored in a database for further usage (red cylinder). The database can flexibly be queried according to user interests (curved red arrows). B, C) Verification (B) and performance benchmarking (C) as two exemplary types of data usage. B) Average activity level (firing rate) in each of the 8 neuronal populations 2/3E, …, 6I depicted in panel A, obtained from simulations of the model on four different GPU platforms (see labels at horizontal axis in panel C). C) Real time factor (ratio between wall-clock time $T_{\text{wall}}$ and simulated biological time $T_{\text{model}}=10\,\text{s}$) for four different GPU computing platforms. Error bars (red) in B) and C) depict standard deviations across ten different model realizations (random-number generator [RNG] seeds) and simulation runs for each platform (error bars are partly too small to be visible).
Hydrology use case
@@
The simulation of the hydrologic cycle is fundamental in environmental modeling. The movement and storage of water in the terrestrial system includes diverse processes. Key processes are the infiltration of precipitation into the subsurface, the storage of water in soils, and the removal of water from the soil via plant transpiration and subsurface runoff. A reliable estimation of water fluxes and storages is relevant for a wide range of sectors, like drinking water supply, agriculture, forestry, energy production, and transportation. Hydrologic models simulate hydrologic state variables and fluxes and their interaction. These models are created for different purposes according to the needs of specific sectors outlined above. These range from infrastructure planning for drinking water supply to drought quantification for agriculture and to climate projections for future water management. Hydrologic models represent fundamental processes like snow accumulation and melting, soil infiltration, evaporation and plant transpiration, surface and subsurface runoff, and river routing. Typical output variables are river discharge, evapotranspiration and soil moisture among others. Input variables are precipitation and air temperature. There are a wide range of computational workflows given the diverse purposes that hydrologic models are used for. Some of the most complex workflows are related to climate change projections where input variables are taken from different sources like climate model ensembles[46]. Currently, metadata tracking in such workflows is often very limited and does not contain the comprehensive settings that are necessary to execute these workflows.